Sydney Ltd provides its employees with long service leave entitlements of 13 weeks of paid leave for every 10 years of continuous service. The company recognises a provision for long service leave using the projected unit credit approach as required by AASB 119. No employees have become entitled to long service leave as the company has only been operating for 6 years.

  • The number of employees working for the company is 80
  • The employees have worked for the company for 3 years
  • The percentage of employees that are expected to become entitled to long service leave is 70%
  • The average annual salary of the employees is $75,000
  • The yield on relevant government corporate bonds is 8% per annum
  • Estimated average increase in wages for the next 10 years is expected to be 5% per annum

Required

Write in the box below the balance of the Provision for Long Service Leave account for Sydney Ltd as at 30 June 20X2, measured in accordance with the requirements of AASB 119. Do not include any dollar signs or spaces when you write your answer.
Answer #1

Since Leave Entitlements is only when an employee contioues for 10 years in the Sydney Ltd. All the employees are working 3 years in company and there are 7 more years to get Leave benefits but only 70% employees continue to work and entitled to Lng Service Leave . Therefore ,

Projected Salary = 75000 * (1+.05)^(10-3)  = 105532 $

Now , Earned Long Service Leave earned will be for 3 years which is worked till now

105532 * (13/52) * 3/10 = 7914.93 $

Now , as this will be discounted @ 8% of govt Bonds rate

7914.93 / (1+.08)^(10-3) = 4618.29 $

Now , the probability for such employees who will be entitled are only 70%

Therefore , it will be 4618.29 * 70% = 3232.80 $